How many terms of the sequence tn = 130 - Sn are included for the sum of the series to equal 1690?



1
Expert's answer
2021-02-18T13:48:54-0500
"t_1=130-S_1, S_1=t_1"

"2t_1=130=>t_1=65"

"S_2=65+t_2, t_2=130-S_2=130-(65+t_2)"

"t_2=\dfrac{1}{2}(65)"


"S_3=65+\dfrac{1}{2}(65)+t_3,"

"t_3=130-S_2=130-(65+\dfrac{1}{2}(65)+t_3 )"

"t_3=(\dfrac{1}{2})^2(65)"

"t_1=65, q=\dfrac{1}{2}"


"t_n=t_1q^{n-1}"

"S_n=\displaystyle\sum_{i=1}^nt_i=\displaystyle\sum_{i=1}^n65\big(\dfrac{1}{2}\big)^n"

"0<\dfrac{1}{2}<1, S=\displaystyle\sum_{i=1}^{\infin}t_i=\dfrac{1}{1-\dfrac{1}{2}}=2""S_n=\displaystyle\sum_{i=1}^n65\big(\dfrac{1}{2}\big)^n<65(2)=130"


The sum of the series is not equal 1690 for "n\in\N."


Therefore there are no solutions.